Techuxia Corporation worked on four jobs during October: Job A256, Job A257, Job A258, and Job A260. At the end of October, the job cost sheets for these jobs contained the following data:

Job A256 Job A257 Job A258 Job A260
Beginning balance $ 1,020 $ 510 $ 0 $ 0
Charged to the jobs during October:
Direct materials $ 2,800 $ 3,990 $ 1,470 $ 3,630
Direct labor $ 1,170 $ 850 $ 510 $ 360
Manufacturing overhead applied $ 1,878 $ 1,900 $ 2,660 $ 530
Units completed 202 0 116 0
Units in process at the end of October 0 301 0 269
Units sold during October 151 0 60 0

Jobs A256 and A258 were completed during October. The other two jobs had not yet been completed at the end of October. There was no finished goods inventory on October 1.

In October, overhead was overapplied by $1,300. The company adjusts its cost of goods sold every month for the amount of the underapplied or overapplied overhead.

1. Using the direct method, what is the cost of goods sold for October?

2.  What is the total value of the finished goods inventory at the end of October?

3.  What is the total value of the work in process inventory at the end of October?
Answer #1

Job Cost sheet

Job A256 Job A257 Job A258 Job A260
Beginning balance 1020 510 0 0
Cost added
Direct material 2800 3990 1470 3630
Direct labor 1170 850 510 360
Applied overhead 1878 1900 2660 530
Total product cost 6868 7250 4640 4520
Units 202 116
Unit product cost 34 40

1) Cost of goods sold = (151*34+60*40)-1300 = $6234

2) Ending finished goods inventory = (51*34+56*40) = $3974

3) Total work in process = 7250+4520 = $11770